the length of a rectangle is 6 ft longer than twice the width. if the perimeter is 66 feet, find the length and the width of the rectangle
Answer by FrankM(1040)

L = 2W+6 and of course W = W
so length + width = 2W+6+W = 33 feet
3W+6 = 33 feet
3W = 27 feet
W=9 feet.
And L = 24 feet
These add to 33 for full perimeter of 66.